Ching leaves San Fransisco and travels toward Los Angeles at 50mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60mi/h. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
Answer by checkley71(8403)

50t=60(t-1)
50t=60t-60
60t-50t=60
10t=60
t=60/10
t=6 hours after Ching leaves they will meet.
50*6=300 miles is the distance until they will meet. Before they get to LA.